An RL circuit is connected to a 230 V, 50 Hz supply with a rating of 500 VA. If the resistor is 20 Ω, calculate the value of the inductive reactance


1
Expert's answer
2021-11-29T00:18:11-0500

The impedance:


"Z^2={X^2+R^2},\\\space\\
X^2=Z^2-R^2."

By Ohm's law:


"Z=V^2/P,\\\space\\
X=\sqrt{(V^2/P)^2-R^2}=104\Omega."